OpenStudy (anonymous):

Hello everyone! :D I am here to help with 6 - 10th grade math Preferably Percent Order of Operation Algebra

OpenStudy (anonymous):

Hey I have an Algebra 2 question would those work?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

What is the equation. Please write it out completely

OpenStudy (anonymous):

Okay so 3(a+c)=d-e solve for c and what is the restriction if any

OpenStudy (anonymous):

3a+3c=d−2.718282

OpenStudy (anonymous):

Define e and d

OpenStudy (anonymous):

But its solve for c

OpenStudy (anonymous):

Yes, but in order to find the exact number/estimation, a given value is needed

OpenStudy (anonymous):

Okay thank you so much.